OpenStudy (anonymous):

6=n/4+5 how do you do this with an explaination

OpenStudy (anonymous):

6 = n / 4 + 5 We want to have n on one side and the numbers ( constants ) on the other. We move the 5 to the other side, by removing 5 from both sides. 6 - 5 = n / 4 + 5 - 5 1 = n/4 Now we multiply both sides by 4. 1*4 = n/4 * 4 4 = n

OpenStudy (anonymous):

thanks!